# Invisible Languages of the LLM Universe

### v1

@@ 3.1 Data Sources and Coverage
We analyze comprehensive data on all 7,613 languages documented in the 25th edition of Ethnologue (Lewis et al.,, 2023). Our analysis integrates two primary dimensions:
Vitality measurement combines:
  • First-language speaker counts from Ethnologue, ranging from single-digit (nearly extinct languages) to hundreds of millions (Hindi, Arabic, Spanish)
  • EGIDS (Expanded Graded Intergenerational Disruption Scale) status ratings from 0 (International) to 10 (Extinct), assessing intergenerational transmission and institutional development
Digitality measurement aggregates presence across:
@@
  • Wikipedia: 64 million articles across 300+ language editions
  • Hugging Face: 114,000 datasets and 447,000 models tagged by language
  • Open Language Archives: 474,000 entries from ELAR, AILLA, DoBeS, and other repositories
Following Simons et al., (2022), we employ automated language identification using fastText and CLD3 classifiers with validation sampling. For each language, we compute normalized scores on each dimension, then apply unsupervised clustering (Gaussian Mixture Models with 3 components) to extract composite vitality and digitality dimensions.
The representation score is computed as:
$\text{Representation}=\text{Digitality}_{\text{normalized}}-\text{Vitality}_{\text{normalized}}$ (1)
Negative scores indicate languages whose vitality exceeds digitality (underrepresented online), positive scores indicate the opposite (overrepresented relative to speaker base).
3.2 Categorical Classification
@@
  • Strongholds (+ Vitality / + Digitality): Both scores above median
  • Digital Echoes (– Vitality / + Digitality): Below-median vitality, above-median digitality
  • Fading Voices (– Vitality / – Digitality): Both scores below median
  • Invisible Giants (+ Vitality / – Digitality): Above-median vitality, below-median digitality
3.3 Geographic and Colonial Analysis
To test whether digitality gaps reflect colonial histories, we coded languages by:
  • Geographic region (Africa, Asia, Europe, Americas, Pacific)
  • Colonial history (colonizer nation and duration, coded from historical databases)
  • Official language status (national, regional, none)
  • Writing system availability and Unicode support
We employ logistic regression predicting Invisible Giant status from colonial history variables, controlling for speaker population and geographic region.
3.4 LLM Training Data Analysis
We analyze language distribution in major LLM training datasets:
  • The Pile (Gao et al.,, 2021): 800GB English-focused corpus
  • mC4 (Xue et al.,, 2021): Multilingual C4 covering 101 languages
  • ROOTS (Laurencçon and the BigScience Workshop,, 2022): BLOOM training data, 46 languages
  • OSCAR (Suárez et al.,, 2019): Common Crawl extractions, 166 languages
For each dataset, we quantify token counts by language and correlate with vitality scores to assess whether training data allocation reflects demographic reality or digital bias.
@@ 4 Results
4.1 The Four-Category Framework
Applying the invisible information framework to all 7,613 documented human languages reveals a stark reality: linguistic representation in digital ecosystems follows systematic patterns that diverge dramatically from demographic vitality. Figure 1 visualizes the vitality-digitality matrix, with each language plotted according to its real-world demographic strength (x-axis) and digital presence across web pages, Wikipedia, datasets, models, and archives (y-axis). The representation score – computed as digitality minus vitality – encodes the gap, with blue indicating digital overrepresentation and red indicating systematic underrepresentation.
Four distinct clusters emerge, each revealing different dynamics of visibility and erasure:
@@ 4.1 The Four-Category Framework
Strongholds (+ Vitality / + Digitality): Roughly one-third of languages occupy the upper-right quadrant, exhibiting both demographic robustness and rich digital representation. These languages form a plume rising above both zero reference lines, their speaker bases, institutional support, and digital footprints mutually reinforcing. This category includes not only global lingua francas but also regionally dominant languages that have successfully translated offline vitality into digital presence. From an LLM perspective, these languages already power most training corpora and anchor multilingual benchmarks. The policy implication is maintenance rather than intervention – keeping data current while redirecting capacity-building resources toward less fortunate quadrants.
@@
Digital Echoes (– Vitality / + Digitality): Only 6% of languages appear in the upper-left quadrant – languages whose online presence exceeds their dwindling speaker communities. Often characterized by historical prestige, liturgical use, or active diaspora networks, these languages demonstrate that digital corpora can outlive real-world vitality. They appear as a thin blue veil above the vitality origin, illustrating how documentation and archiving can preserve linguistic material even as living use declines. For researchers, Digital Echoes serve as cautionary tales: high digital metrics do not guarantee linguistic health. For policymakers, they represent potential leverage points – using strong digital assets to support revitalization before community fluency erodes further.
@@
Fading Voices (– Vitality / – Digitality): The densest cluster, approximately 2,700 languages (36%), hugs the origin and extends into the lower-left. These languages suffer a double deficit: small, vulnerable speech communities and vanishingly small digital footprints. They form a pale cloud of nearly colorless points, underscoring that when both dimensions are scarce, even the representation gap provides minimal signal. From an LLM perspective, they are effectively invisible. From a linguistic diversity standpoint, they constitute the front lines of loss – without urgent fieldwork and community-led documentation, these voices will fade beyond empirical recovery.
@@
Invisible Giants (+ Vitality / – Digitality): Approximately 2,000 languages (27%) occupy the lower-right quadrant – the critical focus of our analysis. These languages manifest as vivid red points: their vitality extends rightward while digitality hovers below the horizontal reference line, producing the largest positive ”need for action” gap on the color scale. These are languages with millions of active speakers but scant digital representation. The paradox is profound: robust demographic presence coexists with near-total digital absence.
The Invisible Giants category reveals the core thesis of this paper: digital exclusion is not simply a function of endangerment or small speaker populations. Languages can be vibrant, widely spoken, and intergenerationally transmitted while remaining systematically invisible in the digital ecosystems that increasingly define authoritative knowledge and power contemporary AI systems.
@@ 4.2 Theoretical Interpretation: Why Invisible Giants Exist
The existence of Invisible Giants – languages with millions of speakers yet minimal digital presence – cannot be explained by technical resource constraints alone. If digitality simply reflected speaker numbers, the vitality-digitality plot would approximate a diagonal line. Instead, we observe systematic deviation: certain languages punch far above their demographic weight digitally, while others with larger populations remain nearly absent.
This pattern aligns precisely with postcolonial theory’s predictions. Said, (1978) established how colonial powers constructed hierarchies positioning non-Western languages as inferior, requiring civilizing intervention. Phillipson, (1992) documented how British colonial education policies systematically suppressed vernacular languages, privileging English and select indigenous lingua francas for administrative convenience. These hierarchies did not disappear with formal decolonization – they became embedded in digital infrastructure decisions.
The infrastructure for digitality – Unicode encoding, keyboard layouts, spell-checkers, search algorithms, content moderation systems – was built primarily by and for Western languages. When platforms assess ”notability” for Wikipedia articles, ”quality” for search rankings, or ”representativeness” for training data, they apply standards developed in English-speaking contexts. As Noble, (2018) demonstrated for search engines, these ostensibly neutral technical systems encode cultural assumptions that systematically disadvantage marginalized communities.
The concept of epistemic injustice (Fricker,, 2007) illuminates the mechanism. Digital systems commit hermeneutical injustice by failing to provide linguistic resources needed for marginalized communities to interpret and share their experiences. When an Invisible Giant language lacks digital presence, speakers cannot access AI tools in their language, cannot find online educational resources, cannot participate in digital commerce on equal footing. The system renders their knowledge production illegible, positioning them as passive consumers of content produced in dominant languages rather than active knowledge creators.
Critically, the Invisible Giants category demonstrates that this exclusion is structural, not natural. These languages are not ”low-resource” due to inherent deficiency – they are systematically under-resourced by institutions that chose to invest in some languages but not others. The framing matters: ”low-resource” naturalizes scarcity, while ”systematically under-resourced” centers the political decisions that produced it (Dotan et al.,, 2023).
@@ 4.3 Implications for Large Language Models
Performance inequality: Models trained predominantly on Stronghold languages perform dramatically worse on Invisible Giants (Joshi et al.,, 2020). This performance gap is not merely inconvenient – it determines who can access AI-powered education, commerce, healthcare, and governance. As LLMs become infrastructure, linguistic exclusion becomes infrastructural inequality.
Knowledge erasure: LLMs encode not just linguistic patterns but conceptual structures, cultural knowledge, and ways of knowing embedded in training text. When training data excludes Invisible Giants, models cannot represent the knowledge systems, historical narratives, and lived experiences of communities speaking those languages. As Bender et al., (2021) argue, this constitutes epistemic violence at scale – billions of parameters encoding the message that certain languages and knowledge don’t matter.
@@
The geographic map in Figure 2 visualizes these patterns spatially. Africa appears predominantly red, indicating languages with higher vitality than digitality. Sub-Saharan African languages, despite centuries of oral tradition maintaining rich knowledge systems, remain systematically absent from digital archives (Nekoto et al.,, 2020). Meanwhile, Europe appears predominantly blue – even small European languages maintain digital presence exceeding their speaker populations, a legacy of institutional investment in digital infrastructure.
This is not coincidence but the continuation of colonial-era hierarchies through algorithmic means – what Mohamed et al., (2020) term ”AI colonialism.” Just as colonial powers extracted material resources while positioning colonized peoples as lacking civilization, contemporary AI development extracts training data (often from Global South digital labor) while positioning non-Western languages as ”low-resource” requiring benevolent intervention.
@@ 4.4 Validation Through Existing Research
Our empirical findings align with and extend prior fragmented observations. Kornai, (2013) established that only 5% of languages will achieve meaningful digital vitality, identifying digital extinction as independent from offline endangerment. Our four-category framework provides the theoretical structure Kornai’s work anticipated but did not fully develop: Digital Echoes exemplify his observation that archived languages can be digitally present without living use; Invisible Giants demonstrate his insight that lack of Wikipedia presence predicts inability to digitally ascend.
Simons et al., (2022) measured digital language support across 143 platforms, finding only 33 languages (0.4%) achieve full capabilities. Their methodology focused on technical infrastructure (encoding, fonts, keyboards); our framework adds the vitality dimension, revealing that infrastructure availability does not guarantee usage. Some Stronghold languages have infrastructure and active communities; Invisible Giants may have infrastructure but minimal content; Fading Voices lack both.
The pattern of 27% Invisible Giants validates warnings from Bender et al., (2021) that LLM training data scarcity for most languages is not incidental but structural. Their ”stochastic parrots” critique – that LLMs encode biases at scale without understanding – becomes more urgent when we quantify precisely which 27% of demographically robust languages are systematically excluded from training.
Crucially, our framework reveals what prior work could not: the magnitude of the representation gap. It is one thing to observe that some languages lack digital resources; it is another to demonstrate that approximately 2,000 languages with millions of speakers score in the highest ”need for action” category. This quantification transforms the problem from anecdotal concern to measurable crisis requiring systematic intervention.
@@ 5.1 Digital-Epistemic Injustice as Structural Phenomenon
Our findings demonstrate that the vitality-digitality gap is not a technical accident but a structural phenomenon reflecting power relations. The systematic concentration of Invisible Giants – 27% of all languages, representing roughly 2,000 languages with millions of speakers yet minimal digital presence – reveals that digital exclusion operates independently of demographic strength.
@@
This constitutes what we term digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of both digital infrastructure and epistemic authority. This dual exclusion operates at unprecedented scale in LLM development. When billions of parameters encode patterns from training data that systematically excludes these 2,000 languages, the resulting models embed the message that certain ways of knowing, certain knowledge systems, and certain communities don’t matter.
@@
The mechanism operates through self-reinforcing feedback loops. Initial digital inequality leads to training data scarcity, producing poor LLM performance, discouraging digital language use, reinforcing data scarcity. Meanwhile, Stronghold languages – already enjoying institutional support and digital infrastructure – attract platform investment, NLP research funding, and commercial applications, widening gaps. Without intervention, these dynamics accelerate language shift as younger generations conclude that economic participation requires adopting dominant languages.
@@
Critically, this is not inevitable technological progress but the reproduction of colonial hierarchies through algorithmic means. The concentration of Strongholds in Europe (where even minority languages maintain robust digital presence) while Invisible Giants concentrate in formerly colonized regions reveals historical continuities. Colonial language policies positioned indigenous languages as inferior, requiring civilizing intervention; contemporary ”low-resource language” framing positions them as deficient, requiring benevolent development assistance. Both framings naturalize inequality while obscuring its political origins.
5.2 Theoretical Contributions
First, we advance epistemic injustice theory by connecting it to digital infrastructure. Fricker, (2007)’s framework focused on face-to-face testimonial exchanges and localized hermeneutical resources. Helm et al., (2024) extended epistemic injustice to language modeling but treated digitality as given. Our analysis reveals that who gets digitized is itself an exercise of power that precedes and enables all subsequent algorithmic processing.
The decision to invest in Unicode encoding, keyboard development, spell-checkers, and training corpora for some languages but not others determines which communities can practice hermeneutical justice – using linguistic resources to interpret experiences in ways that AI systems recognize as valid. When Invisible Giant languages lack these resources, speakers cannot access AI-powered education, healthcare information, or economic opportunities in their languages. The system denies them epistemic standing, positioning them as consumers of knowledge produced elsewhere rather than legitimate knowledge creators.
Second, we demonstrate how postcolonial hierarchies persist in algorithmic systems. While Said, (1978) analyzed textual representations and Mohamed et al., (2020) documented extractive labor practices, we show continuities in linguistic hierarchies themselves. The vitality-digitality gap reveals how colonial language policies – suppressing vernaculars, imposing European languages, privileging select lingua francas – established hierarchies that LLM training data allocation now reproduces computationally.
Contemporary AI development exhibits what we term algorithmic colonialism: extracting digital labor from Global South communities while systematically excluding their languages from AI benefits, then marketing AI tools back to those communities as development interventions. This mirrors classic colonial resource extraction while adding a distinctly epistemic dimension – not merely extracting material resources but positioning entire knowledge systems as lacking, requiring external validation and technological uplift.
Third, we problematize the ”low-resource language” framing ubiquitous in NLP. This terminology positions underrepresented languages as inherently deficient – they lack resources. Our framework reveals that resource scarcity is a political product, not natural fact. The 2,000 Invisible Giants are not inherently low-resource; they are systematically under-resourced by institutions that chose to invest in some languages while neglecting others.
Reframing from ”low-resource” to ”systematically under-resourced” or ”digitally marginalized” centers structural forces rather than naturalizing inequality. This shift has implications beyond terminology – it redirects attention from technical data collection challenges to political questions about whose languages receive institutional investment, whose knowledge systems AI encodes, and who benefits from technological development.

### v2

@@ 3.1 Data Sources and Coverage
We analyze a comprehensive dataset encompassing all 7,613 languages documented in the 25th edition of Ethnologue (eberhard2015ethnologue). The analysis integrates two primary dimensions.
Vitality is operationalized by combining:
  • First-language speaker counts from Ethnologue, which range from single-digit values (for languages on the verge of extinction) to several hundred million speakers (e.g., Hindi, Arabic, Spanish)
  • EGIDS (Expanded Graded Intergenerational Disruption Scale) status ratings, spanning from 0 (International) to 10 (Extinct), which evaluate the degree of intergenerational transmission and the extent of institutional support and development
Digitality is quantified by aggregating a language’s presence across multiple large-scale digital resources:
@@
  • Wikipedia: 64 million articles distributed across more than 300 language editions
  • Hugging Face: 114,000 datasets and 447,000 models annotated with language tags
  • Open Language Archives Community (OLAC): 474,000 catalogued language resource entries
Following simons2022assessing, we implement automated language identification using the fastText and CLD3 classifiers, complemented by validation sampling. For each language, we calculate normalized prevalence scores across each of the aforementioned corpora and subsequently perform principal component analysis–based dimensionality reduction to derive composite indices of linguistic vitality and digital presence.
The degree of visibility of a language on digital platforms is quantified by its representation score, which is defined as:
$\text{Representation}=\text{Digitality}_{\text{normalized}}-\text{Vitality}_{\text{normalized}}\,.$ (1)
Negative values of this score identify languages whose sociolinguistic vitality exceeds their level of digital presence, indicating underrepresentation in online environments. Conversely, positive values correspond to languages that are overrepresented in the digital sphere relative to their speaker base.
@@
  • Strongholds (+ Vitality / + Digitality): Entities characterized by both vitality and digitality scores exceeding the sample median.
  • Digital Echoes (– Vitality / + Digitality): Entities exhibiting below-median vitality combined with above-median digitality.
  • Fading Voices (– Vitality / – Digitality): Entities for which both vitality and digitality scores fall below the median.
  • Invisible Giants (+ Vitality / – Digitality): Entities displaying above-median vitality in conjunction with below-median digitality.
@@ 4 Results
4.1 The Four-Category Framework
Application of the information visibility framework to all 7,613 documented human languages reveals a pronounced systemic disparity: patterns of linguistic representation within digital ecosystems diverge substantially from patterns of demographic vitality. Figure 1 depicts the resulting vitality–digitality matrix, in which each language is positioned according to its real-world demographic strength (x-axis) and its aggregated digital presence across web pages, Wikipedia, datasets, computational models, and archival resources (y-axis). The representation score – defined as the difference between digitality and vitality – quantifies this discrepancy, with blue values indicating digital over-representation and red values indicating systematic under-representation. Four distinct clusters emerge, each revealing different dynamics of visibility and erasure. The geographic spread of these languages can be seen 2, highlighting the stark colonial legacies of invisibility in parts of Africa, Asia, and Oceania.
@@ 4.1 The Four-Category Framework
Strongholds (+ Vitality / + Digitality): Approximately one-third of the languages in the sample fall within the upper-right quadrant, characterized by both high demographic vitality and substantial digital representation (see Figure 3). These languages appear as a plume extending above both zero-reference axes, with speaker populations, institutional backing, and online presence functioning as mutually reinforcing factors. This group comprises not only global lingua francas but also regionally dominant languages that have effectively converted offline sociolinguistic strength into robust digital visibility. From the standpoint of large language models (LLMs), these languages already constitute the bulk of training corpora and serve as primary reference points for multilingual evaluation benchmarks. The principal policy implication is thus one of sustained maintenance rather than active remediation, ensuring that available data remain current while reallocating capacity-building efforts and resources to languages situated in more vulnerable quadrants.
@@
Digital Echoes (– Vitality / + Digitality): Only 6% of languages are located in the upper-left quadrant, representing cases in which digital presence exceeds that of shrinking speaker communities (see Figure 4). Typically associated with historical prestige, liturgical or ritual functions, and/or active diaspora networks, these languages illustrate that digitally available corpora can persist beyond the decline of face-to-face communicative practices. In the visualization, they appear as a thin blue band above the vitality origin, emphasizing how documentation, standardization, and archiving can preserve linguistic material even as inter-generational transmission and everyday use diminish.
From a research perspective, Digital Echoes underscore a critical methodological limitation: elevated digital indicators (e.g., corpus size, number of online documents, or web traffic) do not equate to robust linguistic health. This disconnect arises because most widely used digital metrics are usage-agnostic with respect to (i) the age structure and size of the speaker population, (ii) the domains of actual use (home, school, government, workplace, etc.), (iii) transmission to younger generations, and (iv) speakers’ language attitudes, identity functions, and community control over language resources. A language may thus exhibit substantial digital output generated by a small, aging, or highly specialized group (such as clergy, scholars, or heritage enthusiasts), or by nonnative learners and institutions, while the number of fluent, everyday speakers continues to contract. Conversely, automated or large-scale digitization (e.g., of historical texts) can inflate digital volume without indicating current communicative vitality or community-driven use. For policymakers, these languages constitute strategic leverage points: existing digital infrastructures – including corpora, standardized orthographies, pedagogical materials, and online communities – can be mobilized in support of revitalization policies, such as the development of curricular resources, digital learning tools, and community media, prior to further erosion of any remaining active fluency and functional domains of use.
@@
Fading Voices (– Vitality / – Digitality): The densest cluster, comprising approximately 2,700 languages (36%), is concentrated near the origin of the space and extends into the lower-left quadrant (see Figure 5). These languages exhibit a dual deficit: they are characterized by small, vulnerable speech communities and extremely limited digital presence. In the visualization, they appear as a diffuse cloud of nearly colorless points, indicating that when both dimensions are severely constrained, even the representation-gap metric yields minimal informative signal. From the perspective of large language models (LLMs), these languages are effectively absent from the training landscape. From the perspective of linguistic diversity and language endangerment, however, they represent the bleeding edge of potential loss: without rapid, systematic field documentation and community-driven efforts to record and maintain linguistic practices, these languages are at high risk of disappearing beyond the possibility of empirical reconstruction.
@@
Invisible Giants (+ Vitality / – Digitality): Approximately 2,000 languages (27%) are located in the lower-right quadrant, which constitutes the primary area of analytical concern in this study (see Figure 6). These languages appear as intensely red points: their vitality extends toward the positive end of the horizontal axis, while their digitality remains below the horizontal reference line, yielding the largest positive ”need for action” differential on the color scale. These are languages with millions of active speakers but minimal or negligible digital representation. The paradox is striking: substantial demographic presence coexists with an almost complete absence from digital domains.
The Invisible Giants category encapsulates the central argument of this work: digital exclusion cannot be reduced to linguistic endangerment or to small speaker populations. Languages may be socio-linguistically robust, widely used, demographically significant, and transmitted across generations, while remaining systematically marginalized or absent within the digital ecosystems that increasingly structure authoritative knowledge production and shape the operation of contemporary AI systems.
@@ 4.2 Theoretical Interpretation: Why Invisible Giants Exist
The existence of Invisible Giants – languages with millions of speakers yet minimal digital presence – cannot be explained by technical resource constraints alone. If digitality simply reflected speaker numbers, the vitality-digitality plot would approximate a diagonal line. Instead, we observe systematic deviation: certain languages punch far above their demographic weight digitally, while others with larger populations remain nearly absent. This pattern aligns precisely with postcolonial theory’s predictions. said1977orientalism established how colonial powers constructed hierarchies positioning non-Western languages as inferior, requiring civilizing intervention. These hierarchies did not disappear with formal decolonization – they became embedded in digital infrastructure decisions.
The infrastructure for digitality – Unicode encoding, keyboard layouts, spell-checkers, search algorithms, content moderation systems – was built primarily by and for Western languages. When platforms assess ‘notability’ for Wikipedia articles, ‘quality’ for search rankings, or ‘representativeness’ for training data, they apply standards developed in English-speaking contexts. For search engines, these ostensibly neutral technical systems encode cultural assumptions that systematically disadvantage marginalized communities. The concept of epistemic injustice sheds light on the mechanism. Digital systems commit hermeneutical injustice by failing to provide linguistic resources needed for marginalized communities to interpret and share their experiences. When an Invisible Giant language lacks digital presence, it is hard for the speakers access AI tools in their language, find online educational resources or just participate in digital commerce on equal footing. The system renders their knowledge production illegible, positioning them as passive consumers of content produced in dominant languages rather than active knowledge creators.
Critically, the Invisible Giants category demonstrates that this exclusion is structural, not natural. These languages are not ”low-resource” due to inherent deficiency – they are systematically under-resourced by institutions that chose to invest in some languages but not others. The framing matters: ‘low-resource’ naturalizes scarcity, while ‘systematically under-resourced’ centers the political decisions that produced it (adebara2022towards).
@@ 4.3 Implications for Large Language Models
  • Performance inequality: Models trained predominantly on Stronghold languages perform dramatically worse on Invisible Giants. This performance gap is not merely inconvenient, but also determines who can access AI-powered education, commerce, healthcare, and governance. As LLMs become infrastructure, linguistic exclusion becomes infrastructural inequality.
  • Knowledge erasure: LLMs encode not just linguistic patterns but conceptual structures, cultural knowledge, and ways of knowing embedded in training text. When training data excludes Invisible Giants, models cannot represent the knowledge systems, historical narratives, and lived experiences of communities speaking those languages. This constitutes epistemic violence at scale – billions of parameters encoding the message that certain languages and knowledge don’t matter.
@@
The geographic map in Figure 2 provides a spatial representation of these patterns. Much of Africa appears in red, indicating languages whose sociocultural vitality exceeds their degree of digital representation. Many Sub-Saharan African languages, despite being sustained for centuries through robust oral traditions and complex knowledge systems, remain structurally underrepresented or entirely absent in digital repositories. By contrast, much of Europe appears in blue, reflecting that even relatively small European languages exhibit a level of digital presence that surpasses what would be expected from their speaker populations alone, largely due to long-term institutional investment in digital infrastructures. This distribution does not constitute a random pattern but rather reflects the re-articulation of colonial-era hierarchies through contemporary algorithmic systems. Analogous to how colonial regimes extracted material resources while representing colonized populations as lacking civilization, current AI development practices frequently rely on training data derived from digital labor concentrated in the Global South, while simultaneously classifying non-Western languages as ‘low-resource’ and framing them as objects of technocratic or philanthropic intervention.
@@ 4.4 Validation Through Existing Research
Our empirical results both corroborate and refine prior, largely fragmentary, observations. kornai2013digital estimated that only 5% of the world’s languages are likely to attain substantive digital vitality and argued that digital extinction proceeds largely independently of offline endangerment trajectories. The four-category framework we propose supplies the theoretical architecture that Kornai’s analysis implicitly called for. In particular, Digital Echoes operationalize his claim that languages preserved in archival resources may exhibit a form of digital presence without corresponding contemporary communicative use, while Invisible Giants instantiate his finding that absence from Wikipedia is a strong predictor of a language’s inability to attain higher levels of digital presence.
simons2022assessing conducted a large-scale assessment of digital language support across 143 platforms and found that only 33 languages (0.4%) attained full functional capabilities. While their methodology primarily emphasizes the dimension of technical infrastructure support, our framework introduces an additional vitality perspective, demonstrating that the mere availability of infrastructure does not ensure actual language use. Within our typology, some Stronghold languages exhibit both robust infrastructure and active user communities; Invisible Giants may possess comparable infrastructural provision yet display minimal content production; and Fading Voices are characterized by the absence of both adequate infrastructure and sustained usage.
The distributional pattern characterized as the 27% Invisible Giants empirically substantiates the concerns articulated by bender2021dangers, namely that the paucity of LLM training data for the majority of the world’s languages is not accidental but structurally embedded in current data collection and curation pipelines. The ‘stochastic parrots’ critique – according to which LLMs reproduce and amplify biases at scale without genuine understanding – acquires heightened significance when we quantify which specific 27% of demographically robust languages are systematically marginalized or omitted from training corpora.
Crucially, our framework sheds light on an aspect that prior work has not quantified: the magnitude of the representation gap. It is not only important to document that certain languages lack adequate digital resources, but also to demonstrate that approximately 2000 languages, collectively spoken by millions of individuals, fall into the highest ‘need for action’ category. This quantitative characterization reframes the issue from an anecdotal concern into a measurable crisis that necessitates systematic and large-scale intervention.
@@ 5.1 Digital-Epistemic Injustice as Structural Phenomenon
Our findings demonstrate that the vitality-digitality gap is not a technical accident but a structural phenomenon reflecting power relations. The systematic concentration of Invisible Giants – 27% of all languages, representing roughly 2,000 languages with millions of speakers yet minimal digital presence – reveals that digital exclusion operates independently of demographic strength. This constitutes a kind of digital epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production through denial of both digital infrastructure and epistemic authority. This dual exclusion operates at unprecedented scale in LLM development. When billions of parameters encode patterns from training data that systematically excludes these 2000 languages, the resulting models embed the message that certain ways of knowing, certain knowledge systems, and certain communities don’t matter.
@@
The mechanism operates through self-reinforcing feedback loops. Initial digital inequality leads to training data scarcity, producing poor LLM performance, discouraging digital language use and reinforcing data scarcity. Meanwhile, Stronghold languages – already enjoying institutional support and digital infrastructure – attract platform investment, NLP research funding, and commercial applications, widening gaps. Without intervention, these dynamics accelerate language shift as younger generations conclude that economic participation requires adopting dominant languages.
@@
Critically, this is not inevitable technological progress but the reproduction of colonial hierarchies through algorithmic means. The concentration of Strongholds in Europe (where even minority languages maintain robust digital presence) while Invisible Giants concentrate in formerly colonized regions reveals historical continuities. Colonial language policies positioned indigenous languages as inferior, requiring civilizing intervention; contemporary ‘low-resource language’ framing positions them as deficient, requiring benevolent development assistance. Both framings naturalize inequality while obscuring its political origins.